# Planning and Learning in Risk-Aware Restless Multi-Arm Bandit Problem

### v1

@@ Abstract
In restless multi-arm bandits, a central agent is tasked with optimally distributing limited resources across several bandits (arms), with each arm being a Markov decision process. In this work, we generalize the traditional restless multi-arm bandit problem with a risk-neutral objective by incorporating risk-awareness. We establish indexability conditions for the case of a risk-aware objective and provide a solution based on Whittle index. In addition, we address the learning problem when the true transition probabilities are unknown by proposing a Thompson sampling approach and show that it achieves bounded regret that scales sublinearly with the number of episodes and quadratically with the number of arms. The efficacy of our method in reducing risk exposure in restless multi-arm bandits is illustrated through a set of numerical experiments.
@@ 1 Introduction
The restless multi-arm bandit (RMAB) problem is a class of sequential stochastic control problems for dynamic decision-making under uncertainty. In RMABs, a central agent confronts the challenge of allocating limited resources over time among competing options, which we refer to them as arms, where each is characterized by a Markov decision process (MDP). Such a framework has numerous applications in scheduling problems that appear in machine maintenance Glazebrook et al., (2005); Akbarzadeh and Mahajan, 2019b , healthcare Deo et al., (2013), finance Glazebrook et al., (2013), power management in smart grids Wang et al., (2014); Abad and Iyengar, (2016), opportunistic scheduling in networks Liu and Zhao, (2010); Nino-Mora, (2009); Ouyang et al., (2015); Borkar et al., (2018); Wang et al., (2019), and operator allocation in multi-robot systems Dahiya et al., (2022).
@@
The studies of RMABs have primarily focused on risk-neutral objectives under reward maximization or cost minimization. The assumption of risk neutrality, however, is not always suitable in practice, as the level of risk is an essential aspect to consider in real-world applications Xu et al., (2021); Mate et al., (2021). Such applications encompass various domains, including preventive maintenance Amiri et al., (2018), surgery and medical scheduling in healthcare  He et al., (2019); Najjarbashi and Lim, (2019), financial portfolio management, and production lot-sizing Long et al., (2023), where risk-neutral solutions can be impracticable and lead to risky outcomes. In such circumstances, risk-aware policies accounting for potential risks offer resilient solutions, designed to reduce adverse effects of uncertain outcomes, guaranteeing the allocation policy remains effective under uncertainty Rausand, (2013).
@@
In the traditional RMAB with the risk-neutral objective, the key challenge that prevents applying traditional stochastic control methods is the curse of dimensionality in the state space. As the number of arms increases, the computational complexity of identifying the optimal policy grows exponentially. This poses a formidable obstacle in real-world implementations. Whittle, (1988) introduced a scalable and computationally tractable index policy as a heuristic for RMABs. The Whittle index acts as a priority index that highlights the urgency of selecting an arm. In what follows, we refer to this policy as the Whittle index policy Niño-Mora, (2023).
@@
While the Whittle index policy requires a technical condition, widely known as indexability, to be satisfied, since 1980s, most of the studies in the literature of RMABs implement Whittle index policy. In these problems, either the problem structure is such that the indexability is satisfied Jacko and Villar, (2012); Borkar et al., (2018); Yu et al., (2018); Wang et al., (2019), or sufficient conditions under which the problem is indexable are specified Glazebrook et al., (2006); Niño-Mora, (2007); Akbarzadeh and Mahajan, (2022). The literature suggests that Whittle index policy works well in practice  Glazebrook et al., (2006); Niño-Mora, (2007); Akbarzadeh and Mahajan, (2022); Avrachenkov et al., (2013); Wang et al., (2020).
@@
Risk-aware objectives have been studied extensively for MDPs and RL Le Tallec, (2007); Osogami, (2012); Bäuerle and Rieder, (2014); Chow et al., (2015); Mannor et al., (2016); Xu et al., (2023). To the best of our knowledge, only Mate et al., (2021) have considered risk aware RMABs and do so through numerical experiments by using a specific utility function for a binary state partially-observable MDP.
@@
In contrast with that work, our research provides a comprehensive analysis of risk-awareness across more general state spaces, dynamics, and utility functions. This broader approach allows us to uncover deeper insights into risk-aware decision-making in RMABs. The contribution of our work is fourfold. First, we generalize the traditional RMAB with a risk-neutral objective by incorporating risk-awareness that aims to optimize decision-making with respect to the risk criterion. Second, we then provide conditions under which an arm with a risk-aware objective is indexable, enabling the derivation of the Whittle index policy. Third, we address the learning problem when the true transition probabilities are unknown by proposing a Thompson sampling approach, which samples from posterior distributions over the unknown parameters Osband et al., (2013); Russo et al., (2018).
@@
We derive a regret bound that scales sublinearly in the number of episodes and quadratically with the number of arms. It should be noted that applying a conventional reinforcement learning algorithm to the RMAB in a naive manner is inefficient due to the linear growth of regret in the state space of Markov decision processes, whereas for RMABs, regret grows exponentially with the number of arms Akbarzadeh and Mahajan, (2023). Finally, we numerically illustrate the efficacy of our methodology in reducing risk exposure in RMABs in our experiments in the contexts of machine replacement and patient scheduling applications under both planning and learning setups.
@@ 2.1 Notation
All events occur within a finite horizon setup indexed by $t\in\{0,\ldots,T-1\}=:\mathcal{T}$. Random variables and their realizations are denoted by capital and lowercase letters, respectively; for example, $X_{t}$ and $x_{t}$. We use calligraphic letters to denote the set of all realizations, such as $\mathcal{X}$. Let $X_{a:b}:=\{X_{a},\ldots,X_{b}\}$ represent a collection of the random variables from time $a$ to time $b$, and let $\boldsymbol{X}_{t}=(X^{1}_{t},\ldots,X^{N}_{t})$ represent a collection of the random variable for $N$ processes at time $t$. The probability and the expected value of random events are denoted by $\mathbb{P}(\cdot)$ and $\mathbb{E}(\cdot)$, respectively. Let $\mathbb{I}(\cdot,\cdot)$ be an indicator function for the equality of two terms and ${\boldsymbol{1}}_{x}$ denotes a vector of zeros where only the $x$-th element is one. The notation $(X)^{+}$ represents $\max\{0,X\}$. A function $f$ is called superadditive on partially-ordered sets $\mathcal{X}$ and ${\cal Y}$ if given $x_{1},x_{2}\in\mathcal{X}$ and $y_{1},y_{2}\in{\cal Y}$ where $x_{1}\geq x_{2}$ and $y_{1}\geq y_{2}$, then $f(x_{1},y_{1})-f(x_{1},y_{2})\geq f(x_{2},y_{1})-f(x_{2},y_{2}).$
@@ 2.2 Restless Multi-arm Bandit
A restless multi-Arm bandit (RMAB) problem consists of a set of $N$ independent arms $(\mathcal{X}^{i},\mathcal{A},\allowbreak\{P^{i}(a)\}_{a\in\{0,1\}},r^{i},x^{i}_{0})$, $i\in{\cal N}\coloneqq\{1,\ldots,N\}$. An agent observes the state of all arms and may decide to activate up to $M<N$ of them. Let $\boldsymbol{\mathcal{X}}\coloneqq\prod_{i\in\cal N}\mathcal{X}^{i}$ denote the joint state space and let $\boldsymbol{\mathcal{A}}(M)\coloneqq\bigl{\{}{\boldsymbol{a}}=(a^{1},\ldots,a^{n})\in{\cal A}^{n}:\sum_{i=1}^{N}a^{i}\leq M\bigr{\}}$ denote the action set. The immediate reward realized at time $t$ is $\boldsymbol{r}({\boldsymbol{x}}_{t},{\boldsymbol{a}}_{t})\coloneqq\sum_{i\in{\cal N}}r^{i}(x^{i}_{t},a^{i}_{t})$ when the system is in state ${\boldsymbol{x}}_{t}$ and the agent chooses action ${\boldsymbol{a}}_{t}$. Since the arms are independent, the probability of observing state ${\boldsymbol{x}}_{t+1}$, given the state ${\boldsymbol{x}}_{t}$ and the action ${\boldsymbol{a}}_{t}$ is denoted by ${\boldsymbol{P}}({\boldsymbol{x}}_{t+1}|{\boldsymbol{x}}_{t},{\boldsymbol{a}}_{t}):=\prod_{i\in\cal N}P(x^{i}_{t+1}|x^{i}_{t},a^{i}_{t}).$
@@ 2.3 Planning Problems
Let ${\boldsymbol{\pi}}=({\boldsymbol{\pi}}^{1},\ldots,{\boldsymbol{\pi}}^{N}):\boldsymbol{\mathcal{X}}\times\mathcal{T}\to\boldsymbol{\mathcal{A}}(M)$ denote a time-dependent deterministic policy for the system where ${\boldsymbol{\pi}}^{i}\in\mathcal{X}^{i}\times\mathcal{T}\to{\mathcal{A}}^{i}$ defines the action for arm $i$ in the policy of the system, and let ${\boldsymbol{\Pi}}$ be the set of all such time-dependent deterministic Markov policies.
@@
$\boldsymbol{J}_{{\boldsymbol{x}}_{0}}(\boldsymbol{\pi})$ $:=\sum_{i\in\mathcal{N}}J^{i}_{\boldsymbol{x}_{0}}(\boldsymbol{\pi}^{i}),\text{ with }J^{i}_{\boldsymbol{x}_{0}}(\boldsymbol{\pi}^{i}):=\sum_{t=0}^{T-1}r^{i}\left(X^{i}_{t},A^{i}_{t}\right)\biggm{|}_{{\boldsymbol{X}}_{0}=\boldsymbol{x}_{0}},~{}\forall i\in\mathcal{N}.$
Without loss of generality, let $r^{i}(x,a)\in[0,1/T]$ for any $i\in{\mathcal{N}},x\in{\mathcal{X}}^{i}$, and $a\in{\mathcal{A}}$.
We first describe the classical risk-neutral optimization problem as follows (Whittle,, 1988).
@@
Problem 2.3 is a multi-stage stochastic control problem where the optimal policy is known to be Markovian and presumably obtainable by using dynamic programming (Puterman,, 2014). However, as the cardinality of the state space is $\prod_{i\in{\cal N}}|\mathcal{X}^{i}|$, computing the optimal policy is intractable for large $N$. In Section 3, we describe a well-known heuristic known as Whittle index as a solution to tackle this problem.
As discussed earlier, the assumption of risk-neutrality may not be suitable for various practical applications of RMAB. Thus, we generalize the objective to incorporate risk-sensitivity at the level of the total reward generated by each arm. To this end, we leverage an expected utility formulation (see von Neumann and Morgenstern, (1947)), which is commonly used in the literature of risk-aware decision-making (von Neumann and Morgenstern,, 1947; Fishburn,, 1968; Pratt,, 1978). More specifically, a concave or convex utility function models a risk averse or seeking behavior respectively. Tversky and Kahneman, (1974) further suggest using an S-shaped utility function with inflexion point at a reference value in order to model an attitude that maximizes the chances of reaching such a target.
Problem RRMAB: (Risk-Aware RMAB) Given time horizon $T$, a set of nondecreasing utility functions $U^{i}$, $i\in\mathcal{N}$, and a set of $N$ arms $(\mathcal{X}^{i},\mathcal{A},\{P^{i}(a)\}_{a\in\{0,1\}},r^{i},x^{i}_{0})$, $i\in{\cal N}$, where at most $M<N$ of arms can be activated at a time, find a ${\boldsymbol{\pi}}\in{\boldsymbol{\Pi}}$ that maximizes $\sum_{i\in\mathcal{N}}\mathbb{E}\left[U^{i}\left(J^{i}_{\boldsymbol{x}_{0}}(\boldsymbol{\pi}^{i})\right)\right]$, which we denote in short hand as maximizing $\mathbb{E}\left[\boldsymbol{D}_{{\boldsymbol{x}}_{0}}(\boldsymbol{\pi})\right]$ with
$\boldsymbol{D}_{{\boldsymbol{x}}_{0}}(\boldsymbol{\pi})$ $:=\sum_{i\in\mathcal{N}}D^{i}_{\boldsymbol{x}_{0}}(\boldsymbol{\pi}^{i}),\text{ and }D^{i}_{\boldsymbol{x}_{0}}(\boldsymbol{\pi}^{i}):=U^{i}\left(J^{i}_{\boldsymbol{x}_{0}}(\boldsymbol{\pi}^{i})\right),~{}\forall i\in\mathcal{N}.$
@@
Problem 2.3 highlights risk-awareness for each arm, which aligns with scalarization methods for multi-objective setups Marler and Arora, (2010); Gunantara, (2018). Note that solving the risk-aware problem (Problem 2.3) is more difficult than the risk-neutral problem (Problem  2.3) as already in the case $N=1$, the dynamic programming equation must be written on a state space augmented by one continuous state capturing the cumulative reward so far (Bäuerle and Rieder,, 2014). In section 3, we present how Risk-Aware Whittle index for the risk-aware problem (Problem 2.3) can be obtained.

### v2

@@ Abstract
In restless multi-arm bandits, a central agent is tasked with optimally distributing limited resources across several bandits (arms), with each arm being a Markov decision process. In this work, we generalize the traditional restless multi-arm bandit problem with a risk-neutral objective by incorporating risk-awareness. We establish indexability conditions for the case of a risk-aware objective and provide a solution based on Whittle index. In addition, we address the learning problem when the true transition probabilities are unknown by proposing a Thompson sampling approach and show that it achieves bounded regret that scales sublinearly with the number of episodes and quadratically with the number of arms. The efficacy of our method in reducing risk exposure in restless multi-arm bandits is illustrated through a set of numerical experiments in the contexts of machine replacement and patient scheduling applications under both planning and learning setups.
@@ 1 INTRODUCTION
The restless multi-arm bandit (RMAB) problem is a class of sequential stochastic control problems for dynamic decision-making under uncertainty. In RMABs, a central agent confronts the challenge of allocating limited resources over time among competing options, which we refer to as arms, each characterized by a Markov decision process (MDP). Such a framework has numerous applications in scheduling problems that appear in machine maintenance Glazebrook et al. (2005); Akbarzadeh and Mahajan (2019b), healthcare Deo et al. (2013), finance Glazebrook et al. (2013), power management in smart grids Wang et al. (2014); Abad and Iyengar (2016), opportunistic scheduling in networks Liu and Zhao (2010); Nino-Mora (2009); Ouyang et al. (2015); Borkar et al. (2018); Wang et al. (2019), and operator allocation in multi-robot systems Dahiya et al. (2022).
@@
The studies of RMABs have primarily focused on risk-neutral/expected value objectives under reward maximization or cost minimization. The assumption of risk neutrality, however, is not always suitable in practice, as the level of risk is an essential aspect to consider in real-world applications Xu et al. (2021); Mate et al. (2021). Such applications encompass various domains, including preventive maintenance Amiri et al. (2018), surgery and medical scheduling in healthcare  He et al. (2019); Najjarbashi and Lim (2019), financial portfolio management, and production lot-sizing Long et al. (2023), where risk-neutral solutions can be impracticable and lead to risky outcomes. In such circumstances, risk-aware policies accounting for potential risks offer resilient solutions, designed to reduce adverse effects of uncertain outcomes, guaranteeing the allocation policy remains effective even in adverse conditions Rausand (2013).
@@
In the traditional RMAB with the risk-neutral objective, the key challenge that prevents applying traditional stochastic control methods is the curse of dimensionality due to the state space. As the number of arms increases, the computational complexity of identifying the optimal policy grows exponentially. This poses a formidable obstacle in real-world implementations. Whittle (1988) introduced a scalable and computationally tractable index policy as a heuristic for RMABs. The Whittle index acts as a priority index that highlights the urgency of selecting an arm. In what follows, we refer to this policy as the Whittle index policy Niño-Mora (2023).
@@
While the Whittle index policy requires a technical condition, known as indexability, to be satisfied, most of the studies in the literature of RMABs implement Whittle index policy. In these problems, either the problem structure is such that the indexability is satisfied Jacko and Villar (2012); Borkar et al. (2018); Yu et al. (2018); Wang et al. (2019), or sufficient conditions under which the problem is indexable are specified Glazebrook et al. (2006); Niño-Mora (2007); Akbarzadeh and Mahajan (2022). The literature suggests that Whittle index policy works well in practice  Glazebrook et al. (2006); Niño-Mora (2007); Akbarzadeh and Mahajan (2022); Avrachenkov et al. (2013); Wang et al. (2020).
@@
Risk-aware objectives have been studied extensively for MDPs and RL Le Tallec (2007); Osogami (2012); Bäuerle and Rieder (2014); Chow et al. (2015); Mannor et al. (2016); Jaimungal et al. (2022); Xu et al. (2023). To the best of our knowledge, only Mate et al. (2021) have considered risk aware RMABs and do so through numerical experiments by using a specific utility function for a binary state partially-observable MDP.
@@
In contrast with Mate et al. (2021), our research provides a comprehensive analysis of risk-awareness across more general state spaces, dynamics, and utility functions. This broader approach allows us to uncover deeper insights into risk-aware decision-making in RMABs. The contribution of our work is fourfold. First, we generalize the traditional RMAB with a risk-neutral objective by incorporating risk-awareness that aims to optimize decision-making with respect to the risk criterion. Second, we then provide conditions under which an arm with a risk-aware objective is indexable, enabling the derivation of the Whittle index policy. Third, we address the learning problem when the true transition probabilities are unknown by proposing a Thompson sampling approach, which samples from posterior distributions over the unknown parameters Osband et al. (2013); Russo et al. (2018).
A Recent study, Akbarzadeh and Mahajan (2023), has focused on RMABs with unknown transition dynamics, broadening the scope of applications and solution approaches. Note that papers by Liu et al. (2012); Khezeli and Bitar (2017); Xu et al. (2021) adopt another viewpoint toward restless bandits which is not excatly the same as our problem of interest and it is not based on Whittle index policy. They study a variant of RMABs in which there is a single “best” arm delivering the highest stationary reward; the objective is to learn this arm and pull it indefinitely to maximize long-run return. Although such a strategy is computationally tractable, it differs structurally from our optimization strategy.
@@
We adopts the solution proposed in Akbarzadeh and Mahajan (2023) for our finite-horizon risk-aware setup and derive a regret bound that scales sublinearly in the number of episodes and quadratically with the number of arms. It should be noted that applying a conventional reinforcement learning algorithm to the RMAB in a naive manner is inefficient due to the linear growth of regret in the state space of Markov decision processes, which implies that for RMABs, regret grows exponentially with the number of arms Akbarzadeh and Mahajan (2023). Finally, we numerically illustrate the efficacy of our methodology in reducing risk exposure in RMABs in our experiments in the contexts of machine replacement and patient scheduling applications under both planning and learning setups.
@@ 2.1 Notation
All events occur within a finite horizon setup indexed by $t\in\{0,\ldots,T-1\}=:\mathcal{T}$. Random variables and their realizations are denoted by capital and lowercase letters, respectively; for example, $X_{t}$ and $x_{t}$. We use calligraphic letters to denote the set of all realizations, such as $\mathcal{X}$. Let $X_{a:b}:=\{X_{a},\ldots,X_{b}\}$ represent a collection of the random variables from time $a$ to time $b$, and let $\boldsymbol{X}_{t}=(X^{1}_{t},\ldots,X^{N}_{t})$ represent a collection of the random variable for $N$ processes at time $t$. The probability and the expected value of random events are denoted by $\mathbb{P}(\cdot)$ and $\mathbb{E}[\cdot]$, respectively. Let $\mathbb{I}(\cdot,\cdot)$ be an indicator function for the equality of two terms and ${\boldsymbol{1}}_{x}$ denotes a vector of zeros where only the $x$-th element is one. The notation $(X)^{+}$ represents $\max\{0,X\}$. A function $f$ is called superadditive on partially-ordered sets $\mathcal{X}$ and ${\cal Y}$ if given $x_{1},x_{2}\in\mathcal{X}$ and $y_{1},y_{2}\in{\cal Y}$ where $x_{1}\geq x_{2}$ and $y_{1}\geq y_{2}$, then $f(x_{1},y_{1})-f(x_{1},y_{2})\geq f(x_{2},y_{1})-f(x_{2},y_{2}).$
@@ 2.2 Restless Multi-arm Bandit
A restless multi-Arm bandit (RMAB) problem consists of a set of $N$ independent arms $(\mathcal{X}^{i},\mathcal{A},\allowbreak\{P^{i}(a)\}_{a\in\{0,1\}},r^{i},x^{i}_{0})$, $i\in{\cal N}\coloneqq\{1,\ldots,N\}$. An agent observes the state of all arms and may decide to activate up to $M\leq N$ of them. Let $\boldsymbol{\mathcal{X}}\coloneqq\prod_{i\in\cal N}\mathcal{X}^{i}$ denote the joint state space and let $\boldsymbol{\mathcal{A}}(M)\coloneqq\bigl{\{}{\boldsymbol{a}}=(a^{1},\ldots,a^{n})\in{\cal A}^{n}:\sum_{i=1}^{N}a^{i}\leq M\bigr{\}}$ denote the action set. The immediate reward realized at time $t$ is $\boldsymbol{r}({\boldsymbol{x}}_{t},{\boldsymbol{a}}_{t})\coloneqq\sum_{i\in{\cal N}}r^{i}(x^{i}_{t},a^{i}_{t})$ when the system is in state ${\boldsymbol{x}}_{t}$ and the agent chooses action ${\boldsymbol{a}}_{t}$. Let $r^{i}(x^{i}_{t},a^{i}_{t})\in[R_{\min},R_{\max}]$ where $R_{\min}$ and $R_{\max}$ are finite. Since the arms are independent, the probability of observing state ${\boldsymbol{x}}_{t+1}$, given the state ${\boldsymbol{x}}_{t}$ and the action ${\boldsymbol{a}}_{t}$ is denoted by ${\boldsymbol{P}}({\boldsymbol{x}}_{t+1}|{\boldsymbol{x}}_{t},{\boldsymbol{a}}_{t}):=\prod_{i\in\cal N}P(x^{i}_{t+1}|x^{i}_{t},a^{i}_{t}).$
@@ 2.3 Planning Problems
Let ${\boldsymbol{\pi}}=({\boldsymbol{\pi}}^{1},\ldots,{\boldsymbol{\pi}}^{N}):\boldsymbol{\mathcal{X}}\times\mathcal{T}\to\boldsymbol{\mathcal{A}}(M)$ denote a time-dependent deterministic policy for the system where ${\boldsymbol{\pi}}^{i}$ defines the action for arm $i$ in the policy of the system, and let ${\boldsymbol{\Pi}}$ be the set of all such time-dependent deterministic Markov policies.
@@
$\boldsymbol{J}_{{\boldsymbol{x}}_{0}}(\boldsymbol{\pi})$ $:=\sum_{i\in\mathcal{N}}\sum_{t=0}^{T-1}r^{i}\left(X^{i}_{t},A^{i}_{t}\right)\biggm{|}_{\boldsymbol{\pi},{\boldsymbol{X}}_{0}=\boldsymbol{x}_{0}}.$
We first describe the classical risk-neutral optimization problem as follows (Whittle, 1988).
@@
Problem 2.3 is a multi-stage stochastic control problem where the optimal policy is known to be Markovian and presumably obtainable by using dynamic programming (Puterman, 2014). However, as the cardinality of the state space is $\prod_{i\in{\cal N}}|\mathcal{X}^{i}|$, computing the optimal policy is intractable for large $N$. In Section 3, we describe a well-known heuristic known as Whittle index as a solution to tackle this problem.
As discussed earlier, the assumption of risk-neutrality may not be suitable for various practical applications of RMAB. Thus, we generalize the objective to incorporate risk-sensitivity at the level of the total reward generated by each arm. To this end, we leverage an expected utility formulation (see von Neumann and Morgenstern (1947)), which is commonly used in the literature of risk-aware decision-making (von Neumann and Morgenstern, 1947; Fishburn, 1968; Pratt, 1978). More specifically, a concave or convex utility function models a risk averse or seeking behavior respectively. Tversky and Kahneman (1974) further suggest using an S-shaped utility function with inflexion point at a reference value in order to model an attitude that maximizes the chances of reaching such a target.
Problem RRMAB: (Risk-Aware RMAB) Given time horizon $T$, a set of nondecreasing utility functions $U^{i}$, $i\in\mathcal{N}$, and a set of $N$ arms $(\mathcal{X}^{i},\mathcal{A},\{P^{i}(a)\}_{a\in\{0,1\}},r^{i},x^{i}_{0})$, $i\in{\cal N}$, where at most $M\leq N$ of arms can be activated at a time, find a ${\boldsymbol{\pi}}\in{\boldsymbol{\Pi}}$ that maximizes $\mathbb{E}\left[\boldsymbol{D}_{{\boldsymbol{x}}_{0}}(\boldsymbol{\pi})\right]$ defined below
$\boldsymbol{D}_{{\boldsymbol{x}}_{0}}(\boldsymbol{\pi})$ $:=\sum_{i\in\mathcal{N}}U^{i}\left(\sum_{t=0}^{T-1}r^{i}\left(X^{i}_{t},A^{i}_{t}\right)\right)\biggm{|}_{\boldsymbol{\pi},{\boldsymbol{X}}_{0}=\boldsymbol{x}_{0}}.$
@@
Problem 2.3 highlights risk-awareness for each arm, which aligns with scalarization methods for multi-objective setups Marler and Arora (2010); Gunantara (2018). Note that solving the risk-aware problem (Problem 2.3) is more difficult than the risk-neutral problem (Problem  2.3) as already in the case $N=1$, the dynamic programming equation must be written on a state space augmented by one continuous state capturing the cumulative reward so far (Bäuerle and Rieder, 2014). In section 3, we present how the Risk-Aware Whittle index for the risk-aware problem (Problem 2.3) can be obtained.